if a2-6a-1=0 then a2+1/a2=?
please give detail answer about this question
 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
: a^2 + 1/a^2 = (a- 1/a)^2 + 2 = (a^2-1)^2/a^2
= (6a/a)^2
= 36
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)